OpenStudy (anonymous):

find the x intercept of a parabola with vertex (3,-2) and y intercept (0,7)

OpenStudy (amistre64):

(y+2) = (x-3)^2 +7 y = (x-3)^2 +5

OpenStudy (amistre64):

1,-6,14 3 +- sqrt(36 - 4(14))/2 ... aint see it happening

OpenStudy (amistre64):

yes I do lol.... its right there

OpenStudy (anonymous):

write it in this form (x1,y1)(x2,y2)

OpenStudy (amistre64):

(y+2) = (x-3)^2 +C y = (x-3)^2 +C -2 C-2 = 7; C =9 ........................................ y = (x-3)^2 +9 y = x^2 -6x +9 +9; ug, scratch that y = x^2 -6x +7 .............................. 6 +- sqrt(36 -28) sqrt(8)...2sqrt(2) ---------------- = 2 (3 + sqrt(2), 0) (3 - sqrt(2), 0) maybe lol

OpenStudy (anonymous):

\[y=a(x-3)^2-2\] and \[7=a(0-3)^2+2\] \[7=9a+2\] \[5=9a\] \[a=\frac{5}{9}\]

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

giving \[y=\frac{5}{9}(x-3)^2-2\]

OpenStudy (anonymous):

in the (x1,y1) and (x2,y2) form please, i need two intercepts

OpenStudy (anonymous):

ok we have the equation for the parabola. finding x- intercepts means setting y = 0 and solve for x.

OpenStudy (amistre64):

x^2 -6x + 7 is good

OpenStudy (anonymous):

in other words love \[0=\frac{5}{9}(x-3)^2-2\]

OpenStudy (amistre64):

i used 0,7 ; 3,-2 ; 6,7 lol

OpenStudy (anonymous):

in the (x1,y1) and (x2,y2) form

OpenStudy (anonymous):

please

OpenStudy (anonymous):

well x-intercept means \[(x_1,0)\]

OpenStudy (amistre64):

already did; (3-sqrt(2),0) and (3+sqrt(2),0)

OpenStudy (anonymous):

what equation did you use?

OpenStudy (amistre64):

x^2 -6x +7

OpenStudy (anonymous):

(y-k)=a(x-h)^2

OpenStudy (amistre64):

(y-7)+9 = (x-3)^2 (y+2) = (x-3)^2

OpenStudy (anonymous):

boy am i dumb. i wrote \[y=a(x-3)^2+2\] when it should have been \[y=a(x-3)^2-2\] and \[7=a(-3)^2-2\] \[7=9a-2\] \[9=9a\] \[a=1\]

OpenStudy (amistre64):

:) its ok lol

OpenStudy (anonymous):

so equation is \[y=(x-3)^2-2\]

OpenStudy (anonymous):

idk how to change it to the two points on the graph, please type it that way

OpenStudy (anonymous):

or if you prefer \[y+2=(x-3)^2\]

OpenStudy (amistre64):

i posted the points 2 times already...

OpenStudy (amistre64):

(3+sqrt(2),0) (3-sqrt(2),0) right?

OpenStudy (anonymous):

zeros at \[3\pm \sqrt{2}\]

OpenStudy (anonymous):

with the parenthesis? and two different points?

OpenStudy (anonymous):

with the parenthesis? and two different points?

OpenStudy (anonymous):

by inspection

OpenStudy (amistre64):

now that 3 and four times lol

OpenStudy (anonymous):

\[(2 -\sqrt{3},0)\] \[(2+\sqrt{3},0)\]

OpenStudy (amistre64):

sat lol; that was backwards

OpenStudy (anonymous):

are the x-intercepts sorry i slowed you up

OpenStudy (anonymous):

oh yes it was! \[(3-\sqrt{2},0)\] \[(3+\sqrt{2},0)\]

OpenStudy (anonymous):

please in the form of (x1,y1) and (x2,y2) idk what (2-sqrt of 3, 0( means...

OpenStudy (anonymous):

means \[x_1=3-\sqrt{2}\] \[y_1=0\]

OpenStudy (amistre64):

it means the the sqrt(2) is an EXACT form of an irrational number that can only be expressed approximately in decimal form; of which we have no way of determining the requirements of since you have given us no parameters to establish them by

OpenStudy (anonymous):

and \[x_2=3+\sqrt{2}\] \[y_2=0\]

OpenStudy (amistre64):

y = 0 exactly at: x = 3-sqrt(2) AND x= 3-sqrt(2) ; otherwise it doesnt...

OpenStudy (amistre64):

lol... make one of those a +sqrt(2)

OpenStudy (anonymous):

whats x1 and y1?

OpenStudy (amistre64):

x1 = 3-sqrt(2) y1 = 0

OpenStudy (anonymous):

its wrong

OpenStudy (amistre64):

x1 = 1.414213562373095048801688724209.... no, its right; you simply havent given us a means to approximate the answer.

OpenStudy (amistre64):

its like saying; what color is the sky? blue. wrong, its a shade of blue that is determined by whatever the criteria is for saying 'blue' in a way that is acceptable...

OpenStudy (amistre64):

how old am I ? the right answer depends on how accurate of an answer you want right?